# Federico Echenique and Farzad Pourbabaee

### v4

@@ 1 Introduction
The diagram on the left in Figure 1 illustrates the situation, albeit in the case with two states ($d=2$) and two agents. Aggregate consumption is fixed at 1 (say 1 million dollars) in every state, represented by the vector $\omega=(1,1)$. Aggregate consumption is state-independent, and thus implies no aggregate risk. The square defined by the origin and $\omega=(1,1)$ is an Edgeworth box that describes all the allocations—all risk-sharing agreements—between the two agents. The Pareto optimal allocations are on the solid curve connecting the origin and $\omega$, while $f=(f_{1},f_{2})$ is a particular Pareto optimal allocation (indicated by the tangency of the agents’ indifference curves at the point $f$, with the usual Edgeworth-box convention that one agent’s coordinate system has been rotated $180^{\circ}$ and its origin coincides with $\omega$). For the allocation $f$, we may consider all the aggregate consumptions that can be decentralized among the two agents so as to obtain higher utility “by $\varepsilon$” than in $f$ [^2] An improvement by $\varepsilon$ means a state-contingent consumption $g$ so that $(1-\varepsilon)g$ is strictly preferred to $f_{i}$. If, for example, preferences are homothetic, this translates into a utility improvement of $\varepsilon/(1-\varepsilon)$%. This set of aggregate bundles consists of all the bundles on the dotted region to the north-east of the dashed curve passing through $\omega$ (a curve that, at $\omega$, has the same slope as the common tangent to agents’ indifference curves at $f$).
[…]
@@ 2.1 Notations and Conventions
If $S$ is a finite set, we denote by $\Delta S=\{\mu:S\mapsto\mathbb{R}_{+}|\sum_{s\in S}\mu(s)=1\}$ the set of all probability measures on $S$. We embed $\Delta S$ as a subset of $\mathbb{R}_{+}^{d}$, and sometimes we refer to this probability simplex by $\Delta_{d}$.
@@
$\mathcal{M}^{r}_{\kappa}\coloneq\left\{\mathsf{P}^{r}_{\kappa}\in\Delta\left(\mathbb{B}(r)\right):\frac{\mathrm{d}\mathsf{P}^{r}_{\kappa}}{\mathrm{d}\mathsf{P}^{r}}(z)\leq\kappa,\,\,\text{a.e.}\,\,z\in\mathbb{B}(r)\right\}\,.$
@@
For two subsets $A,B\subseteq\mathbb{R}^{m}$, we define $\mathsf{dist}(A,B)\coloneq\inf\left\{\lVert a-b\rVert:a\in A,b\in B\right\}$. We refer to the $\delta$-extension of the subset $A$ by $A^{\delta}=\{x:\mathsf{dist}(x,A)<\delta\}$, which coincides with $A+\delta\,\mathbb{B}$.
Given a measurable subset $A\subseteq\mathbb{R}^{m}$, we denote its Euclidean volume by $\mathsf{Vol}(A)$, which is equal to the Lebesgue integral of the indicator function of $A$ relative to the affine hull of $A$. For example, if $A$ is a $(m-1)$-dimensional surface in $\mathbb{R}^{m}$, then $\mathsf{Vol}(A)$ refers to the surface area of $A$, as opposed to its $m$-dimensional volume (which is zero).
@@ 2.2 Preferences and Uncertainty
We consider a setting with uncertainty. [^3] Using the reinterpretation of commodities in Debreu (1959), our results have implications for many other economic environments. For example, for the textbook version of the model which assumes that consumption is in units of physically distinct goods. Let $S$ be a finite set of possible states of the world, and $d$ be the number of states of the world; $d\coloneq|S|$.
Consequences are evaluated based on their monetary payoff in $\mathbb{R}$. The set of acts, mappings from $S\to\mathbb{R}$, is denoted by $\mathbb{R}^{d}$, and individual acts are denoted by $f$ and $g$. We endow this set with its natural topology.
@@
    That is, for every $f\in\mathbb{R}^{d}$, the sets $\{g:g\succeq f\}$ and $\{g:f\succeq g\}$ are closed subsets of $\mathbb{R}^{d}$.
@@
Given a preference $\succeq$ and an act $f$, the set $\{g:g\succeq f\}$ is called the upper contour set of $\succeq$ at $f$. We say that a preference $\succeq$ is convex if its upper contour sets are convex, for all acts $f$. Convexity jointly with the weak order property and continuity imply that the strict upper contour set, denoted by $\mathcal{U}_{\succeq}^{(0)}\coloneq\{g:g\succ f\}$, is also convex. We refer to the space of convex preferences by $\mathcal{C}\subset\mathcal{P}$. We interpret convexity in our model as a basic property of risk, or uncertainty aversion (Yaari, 1969).
@@
$\mathcal{U}_{\succeq}^{(\varepsilon)}(f)=\left\{g\in\mathbb{R}^{d}:(1-\varepsilon)g\succ f\right\}\,.$
@@
Approximate optimality is often expressed by means of the $\varepsilon$-maximization of some numerical objective function—a utility function $u$ representing a preference $\succeq$. Observe that when $\succeq$ is homothetic (which is the case for many preferences used in applications), then $u$ may be taken to be homogeneous of degree one. As a consequence, our notion of approximate optimality translates directly into an approximation of utilities. Indeed, we may then write $\mathcal{U}_{\succeq}^{(\varepsilon)}(f)=\left\{g\in\mathbb{R}^{d}:(1-\varepsilon)u(g)>u(f)\right\}$. Then an act $g$ is in $\mathcal{U}_{\succeq}^{(\varepsilon)}(f)$ if and only if it represents an $\varepsilon/(1-\varepsilon)$utility improvement over $f$.
@@ 3.1 Walrasian Equilibrium
Our first result entails the following exercise: Fix a Walrasian equilibrium allocation, and draw a random perturbation from a ball of radius $r$ distributed according to some probability measure $\mathsf{P}^{r}_{\kappa}\in\mathcal{M}^{r}_{\kappa}$. [^5] Obviously the interesting measures $\mathsf{P}^{r}_{\kappa}\in\mathcal{M}^{r}_{\kappa}$ involve $\kappa$ that is independent of $d$, which is the case for the uniform and conditional Gaussian measures. See our remarks in Section 2.1. Consider the event that this random perturbation, when applied to any agent’s equilibrium consumption, results in a contingent consumption that is better by at least $\varepsilon$ than their equilibrium consumption. The resulting event, that at least one agent is made better off, has vanishingly small probability.
@@
Let $\mathcal{E}\colon I\to\mathcal{P}\times\mathbb{R}_{+}^{d}$ be an exchange economy. Suppose that there exists $\tau>0$ such that $\omega_{i}\geq\tau\bm{1}$ for all $i\in I$, and let $f$ be a Walrasian equilibrium allocation. Fix $r>0$ and let $z\sim\mathsf{P}^{r}_{\kappa}\in\mathcal{M}^{r}_{\kappa}$. Then for every $\varepsilon>0$,
@@
The proof of this theorem is in Section 5, as is the proof of the rest of the results we provide in this section. The proof relies on results from the literature on isoperimetric inequalities and the concentration of measure phenomenon, but the proof section attempts to be self-contained. [^6] One simple intuition highlights the connection with the concentration of measures: Consider an economy with a single risk-neutral agent, and assume that all states are equally likely; so the agent’s preferences are represented by $u(f)=\sum_{s}f(s)$. Then an $\varepsilon$-improvement over a constant act $f(s)=1$ (which would be an equilibrium allocation in a Robinson Crusoe economy) requires that $(1-\varepsilon)\sum_{s}z(s)>\varepsilon d$, or that $\frac{1}{d}\sum_{s}z(s)>\varepsilon/(1-\varepsilon)$. In high dimensions, despite the lack of independence ($z$ is drawn from a sphere), the random variable $\frac{1}{d}\sum_{s}z(s)$ is tightly concentrated around 00.
@@ 3.3 Prior Beliefs and Welfare-Improving Trade
Specifically, let $f$ be an act in $\mathbb{R}_{+}^{d}$. The upper contour set of agent $i$ is $\{g:g\succeq_{i}f\}$, and the subjective belief set at $f$ is defined by
$\mathcal{B}_{i}(f)=\left\{\mu\in\Delta S:\mu\cdot g\geq\mu\cdot f\text{ for all }g\succeq_{i}f\right\}\,.$
The set $\mathcal{B}_{i}(f)$ is a convex and closed (hence compact) subset of the $(d-1)$-dimensional probability simplex $\Delta S$ (or $\Delta_{d}$). One may also interpret a vector in $\mathcal{B}_{i}(f)$ as the vector of normalized prices that supports the consumption of the act $f$.
@@
$\mathcal{B}_{i}(f_{i})^{\delta}\coloneq\{\nu\in\Delta S:\inf_{\mu\in\mathcal{B}_{i}(f_{i})}\lVert\nu-\mu\rVert<\delta\}\,.$
@@ 4.1 Behavioral Implications of Theorem 3
We further assume that $\Pi$ has constant width, namely that the distance between any distinct parallel supporting hyperplanes of $\Pi$ (residing on the $(d-1)$-dimensional probability simplex) is constant. This assumption is crucial for our subsequent use of the result of Schramm (1988), but one can imagine a quantitative relaxation of the assumption of constant width and that it would translate into a modified bound on the degree of ambiguity aversion.
@@ 5.1 Brunn-Minkowski Inequality
For two subsets $A,B\subseteq\mathbb{R}^{d}$, their Minkowski sum is defined by $A+B\coloneq\left\{a+b:a\in A,b\in B\right\}$. The Brunn-Minkowski inequality provides a crucial connection between volumes and the Minkowski sum in Euclidean spaces.
@@
We proceed with the proof of our first two theorems. In both cases, using the optimality or the equilibrium property of the allocation, we apply a type of convex separation argument. In the first theorem, the separation is provided by the given equilibrium price. While in the second theorem, the separation argument follows the idea used in the proof of the second welfare theorem. Consequently, as we argue below, the approximate versions of upper contour sets stay at a positive distance from their separating counterpart. An application of Lemma 1 to appropriately chosen subsets implies the probability bounds in Theorems 1 and 2.
We begin by laying down some terminology. For a vector $p\in\mathbb{R}^{d}$ and a constant $b$, we define two half-spaces:
$\begin{gathered}H^{+}(p\,;\,b)\coloneq\left\{x\in\mathbb{R}^{d}:p\,\cdot\,x\geq b\right\}\,,\\ H^{-}(p\,;\,b)\coloneq\left\{x\in\mathbb{R}^{d}:p\,\cdot\,x\leq b\right\}\,,\end{gathered}$
@@
where the last inequality holds because $p\in\mathbb{R}_{+}^{d}$ and $\omega_{i}\geq\tau\bm{1}$. Therefore, since $i\in I$ and $g_{i}\in\mathcal{U}_{i}^{(\varepsilon)}(f_{i})$ were arbitrary, by the above inequality $\mathcal{U}_{i}^{(\varepsilon)}(f_{i})-\{\omega_{i}\}\subseteq H^{+}\left(p;\varepsilon\tau\lVert p\rVert_{1}\right)$ for all $i\in I$. Let us define $\mathcal{Q}\coloneq\bigcup_{i\in I}\left(\mathcal{U}_{i}^{(\varepsilon)}(f_{i})-\{\omega_{i}\}\right)$. [^9] The set $\mathcal{Q}$ was originally used by Debreu and Scarf (1963) to prove core convergence, and by Barman and Echenique (2023) to characterize approximate Walrasian equilibria. Then $\mathcal{Q}\subseteq H^{+}\left(p;\varepsilon\tau\lVert p\rVert_{1}\right)$, so for an arbitrary $r>0$, one has
@@
The equality above follows from (5.4), and the last inequality holds because $\min_{p\neq\bm{0}}\lVert p\rVert_{1}/\lVert p\rVert_{2}=1$, namely the minimum of $\lVert p\rVert_{1}/\lVert p\rVert_{2}$ is achieved on the standard unit basis vectors, and is equal to $1$. Now set $A\coloneq\mathcal{Q}\cap\mathbb{B}(r)$ and $B\coloneq H^{-}(p;0)\cap\mathbb{B}(r)$. By the above inequality $\mathsf{dist}(A,B)\geq\varepsilon\tau$. Since $p$ is a nonzero vector in $\mathbb{R}_{+}^{d}$, the subset $B$ covers at least half of the volume of $\mathbb{B}(r)$. So it must be that $\mathsf{Vol}(A)\leq\mathsf{Vol}(B)$. Therefore, Lemma 1 implies that $\mathsf{Vol}(A)/\mathsf{Vol}(\mathbb{B}(r))\leq\mathrm{e}^{-\varepsilon^{2}\tau^{2}d/8r^{2}}$, namely:
@@
Now suppose $v\in\mathcal{V}^{(\varepsilon)}$. Then, there are $g_{i}\in\mathcal{U}_{i}^{(\varepsilon)}$ for every $i\in I$, such that $v=\sum_{i\in I}g_{i}$ and $(1-\varepsilon)g_{i}\succ_{i}f_{i}$ For each $i\in I$, it holds that $(1-\varepsilon)g_{i}\in\mathcal{U}_{i}^{(0)}$ and hence $(1-\varepsilon)\sum_{i\in I}g_{i}\in\mathcal{V}^{(0)}$. Consequently, one has $p\cdot(1-\varepsilon)v\geq\lVert p\rVert_{1}$. That in turn implies $v\in H^{+}\left(p;\lVert p\rVert_{1}/(1-\varepsilon)\right)$, and thereby $\mathcal{V}^{(\varepsilon)}\subseteq H^{+}\left(p;\lVert p\rVert_{1}/(1-\varepsilon)\right)$. As a result of this set inclusion, for an arbitrary $r>0$, one obtains that
$\begin{gathered}\mathsf{dist}\left(\mathcal{V}^{(\varepsilon)}\cap\mathbb{B}(\bm{1},r)\,,\,H^{-}(p\,;\,\lVert p\rVert_{1})\cap\mathbb{B}(\bm{1},r)\right)\geq\\ \mathsf{dist}\left(H^{+}\left(p\,;\,\frac{\lVert p\rVert_{1}}{1-\varepsilon}\right)\cap\mathbb{B}(\bm{1},r)\,,\,H^{-}\left(p\,;\,\lVert p\rVert_{1}\right)\cap\mathbb{B}(\bm{1},r)\right)\\ =\mathsf{dist}\left(H^{+}\left(p\,;\,\frac{\lVert p\rVert_{1}}{1-\varepsilon}\right)\,,\,H^{-}(p\,;\,\lVert p\rVert_{1})\right)\,,\end{gathered}$

### v5

@@ 1 Introduction
The diagram on the left in Figure 1 illustrates the situation, albeit in the case with two states ($d=2$) and two agents. Aggregate consumption is fixed at 1 (say 1 million dollars) in every state, represented by the vector $\omega=(1,1)$. Aggregate consumption is state-independent, and thus implies no aggregate risk. The square defined by the origin and $\omega=(1,1)$ is an Edgeworth box that describes all the allocations—all risk-sharing agreements—between the two agents. The Pareto optimal allocations are on the solid curve connecting the origin and $\omega$, while $f=(f_{1},f_{2})$ is a particular Pareto optimal allocation (indicated by the tangency of the agents’ indifference curves at the point $f$, with the usual Edgeworth-box convention that one agent’s coordinate system has been rotated $180^{\circ}$ and its origin coincides with $\omega$). For the allocation $f$, we may consider all the aggregate consumptions that can be decentralized among the two agents so as to obtain higher utility “by $\varepsilon$” than in ff. [^2] An improvement by $\varepsilon$ means a state-contingent consumption $g$ so that $(1-\varepsilon)g$ is strictly preferred to $f_{i}$. If, for example, preferences are homothetic, this translates into a utility improvement of $\varepsilon/(1-\varepsilon)$%. This set of aggregate bundles consists of all the bundles on the dotted region to the north-east of the dashed curve passing through $\omega$ (a curve that, at $\omega$, has the same slope as the common tangent to agents’ indifference curves at $f$).
$f(s_{1})$$f(s_{2})$$\omega$1111ff
@@ 2.1 Notations and Conventions
If $S$ is a finite set, we denote by $\Delta S=\{\mu:S\mapsto\mathbb{R}_{+}\mid\sum_{s\in S}\mu(s)=1\}$ the set of all probability measures on $S$. We embed $\Delta S$ as a subset of $\mathbb{R}_{+}^{d}$, and sometimes we refer to this probability simplex by $\Delta_{d}$.
@@
$\mathcal{M}^{r}_{\kappa}\coloneq\left\{\mathsf{P}^{r}_{\kappa}\in\Delta\left(\mathbb{B}(r)\right)\colon\frac{\mathrm{d}\mathsf{P}^{r}_{\kappa}}{\mathrm{d}\mathsf{P}^{r}}(z)\leq\kappa,\,\,\text{a.e.}\,\,z\in\mathbb{B}(r)\right\}\,.$
@@
For two subsets $A,B\subseteq\mathbb{R}^{m}$, we define $\mathsf{dist}(A,B)\coloneq\inf\left\{\lVert a-b\rVert\colon a\in A,b\in B\right\}$. We denote the $\delta$-extension of the subset $A$ by $A^{\delta}=\{x\colon\mathsf{dist}(x,A)<\delta\}$, which coincides with $A+\delta\,\mathbb{B}$.
Given a measurable subset $A\subseteq\mathbb{R}^{m}$, we denote its Euclidean volume by $\mathsf{Vol}(A)$, which is equal to the Lebesgue integral of the indicator function of $A$ relative to the affine hull of $A$. For example, if $A$ is a (m1)(m-1)-dimensional surface in $\mathbb{R}^{m}$, then $\mathsf{Vol}(A)$ refers to the surface area of $A$, as opposed to its $m$-dimensional volume (which is zero).
@@ 2.2 Preferences and Uncertainty
We consider a setting with uncertainty. [^3] Using the reinterpretation of commodities in Debreu (1959), our results have implications for many other economic environments. For example, they apply to the textbook version of the model, which assumes that consumption is in units of physically distinct goods. Let $S$ be a finite set of possible states of the world, and $d$ be the number of states of the world; $d\coloneq|S|$.
Consequences are evaluated based on their monetary payoff in $\mathbb{R}$. The set of acts, mappings from $S\to\mathbb{R}$, is denoted by $\mathbb{R}^{d}$, and individual acts are denoted by $f$ and $g$. We endow this set with its standard topology.
@@
    That is, for every $f\in\mathbb{R}^{d}$, the sets $\{g\colon g\succeq f\}$ and $\{g\colon f\succeq g\}$ are closed subsets of $\mathbb{R}^{d}$.
@@
Given a preference $\succeq$ and an act $f$, the set $\{g\colon g\succeq f\}$ is called the upper contour set of $\succeq$ at $f$. We say that a preference $\succeq$ is convex if its upper contour sets are convex, for all acts $f$. Convexity jointly with the weak order property and continuity imply that the strict upper contour set, denoted by $\mathcal{U}_{\succeq}^{(0)}(f)\coloneq\{g\colon g\succ f\}$, is also convex. We refer to the space of convex preferences by $\mathcal{C}\subset\mathcal{P}$. We interpret convexity in our model as a basic property of risk, or uncertainty aversion (Yaari, 1969).
@@
$\mathcal{U}_{\succeq}^{(\varepsilon)}(f)=\left\{g\in\mathbb{R}^{d}\colon(1-\varepsilon)g\succ f\right\}\,.$
@@
Approximate optimality is often expressed by means of the $\varepsilon$-maximization of some numerical objective function—a utility function $u$ representing a preference $\succeq$. Observe that when $\succeq$ is homothetic (which is the case for many preferences used in applications), then $u$ may be taken to be homogeneous of degree one. As a consequence, our notion of approximate optimality translates directly into an approximation of utilities. Indeed, we may then write $\mathcal{U}_{\succeq}^{(\varepsilon)}(f)=\left\{g\in\mathbb{R}^{d}\colon(1-\varepsilon)u(g)>u(f)\right\}$. Then an act $g$ is in $\mathcal{U}_{\succeq}^{(\varepsilon)}(f)$ if and only if it represents an $\varepsilon/(1-\varepsilon)$% utility improvement over $f$.
@@ 3.1 Walrasian Equilibrium
Our first result entails the following exercise: Fix a Walrasian equilibrium allocation, and draw a random perturbation from a ball of radius $r$ distributed according to some probability measure $\mathsf{P}^{r}_{\kappa}\in\mathcal{M}^{r}_{\kappa}$. [^5] Obviously the interesting measures $\mathsf{P}^{r}_{\kappa}\in\mathcal{M}^{r}_{\kappa}$ involve $\kappa$ that is independent of $d$, which is the case for the uniform and conditional Gaussian measures. See our remarks in Section 2.1. Consider the event that this random perturbation, when applied to any agent’s equilibrium consumption, results in a contingent consumption that is better by at least $\varepsilon$ than their equilibrium consumption. The resulting event, that at least one agent is made better off, has a vanishingly small probability as $d$ increases.
@@
Let $\mathcal{E}\colon I\to\mathcal{P}\times\mathbb{R}_{+}^{d}$ be an exchange economy, $\omega=\left(\omega_{i}\right)_{i\in I}$ be the endowment profile, and $f=\left(f_{i}\right)_{i\in I}$ be a Walrasian equilibrium allocation. Assume that there exists $\tau>0$ such that $\omega_{i}\geq\tau\bm{1}$ and $f_{i}\geq\tau\bm{1}$ for all $i\in I$. Fix $r>0$ and let $z\sim\mathsf{P}^{r}_{\kappa}\in\mathcal{M}^{r}_{\kappa}$. Then for every $\varepsilon>0$,
@@
The proof of this theorem is in Section 5, as is the proof of the rest of the results we provide in this section. The proof relies on results from the literature on isoperimetric inequalities and the concentration of measure phenomenon, but the proof section attempts to be self-contained. [^6] One simple intuition highlights the connection with the concentration of measures: Consider an economy with a single risk-neutral agent, and assume that all states are equally likely; so the agent’s preferences are represented by $u(f)=\sum_{s}f(s)$. Then an $\varepsilon$-improvement over a constant act $f(s)=1$ (which would be an equilibrium allocation in a Robinson Crusoe economy) requires that $(1-\varepsilon)\sum_{s}z(s)>\varepsilon d$, or that $\frac{1}{d}\sum_{s}z(s)>\varepsilon/(1-\varepsilon)$. In high dimensions, despite the lack of independence ($z$ is drawn from a sphere), the random variable $\frac{1}{d}\sum_{s}z(s)$ is tightly concentrated around 0.
@@ 3.3 Prior Beliefs and Welfare-Improving Trade
Specifically, let $f$ be an act in $\mathbb{R}_{+}^{d}$. The upper contour set of agent $i$ is $\{g\colon g\succeq_{i}f\}$, and the subjective belief set at $f$ is defined by
$\mathcal{B}_{i}(f)=\left\{\mu\in\Delta S\colon\mu\cdot g\geq\mu\cdot f\text{ for all }g\succeq_{i}f\right\}\,.$
The set $\mathcal{B}_{i}(f)$ is a convex and closed (hence compact) subset of the (d1)(d-1)-dimensional probability simplex $\Delta S$ (or $\Delta_{d}$). One may also interpret a vector in $\mathcal{B}_{i}(f)$ as the vector of normalized prices that supports the consumption of the act $f$.
@@
$\mathcal{B}_{i}(f_{i})^{\delta}\coloneq\{\nu\in\Delta S\colon\inf_{\mu\in\mathcal{B}_{i}(f_{i})}\lVert\nu-\mu\rVert<\delta\}\,.$
@@ 4.1 Behavioral Implications of Theorem 3
We further assume that $\Pi$ has constant width, namely that the distance between any distinct parallel supporting hyperplanes of $\Pi$ (residing on the (d1)(d-1)-dimensional probability simplex) is constant. This assumption is crucial for our subsequent use of the result of Schramm (1988), but one can imagine a quantitative relaxation of the assumption of constant width and that it would translate into a modified bound on the degree of ambiguity aversion.
@@ 5.1 Brunn-Minkowski Inequality
For two subsets $A,B\subseteq\mathbb{R}^{d}$, their Minkowski sum is defined by $A+B\coloneq\left\{a+b\colon a\in A,b\in B\right\}$. The Brunn-Minkowski inequality provides a crucial connection between volumes and the Minkowski sum in Euclidean spaces.
@@
We proceed with the proof of our first two theorems. In both cases, using the optimality or the equilibrium property of the allocation, we apply a type of convex separation argument. In the first theorem, the separation is provided by the given equilibrium price. In the second theorem, the separation argument follows the ideas used in the standard proofs of the second welfare theorem. Consequently, as we argue below, the approximate versions of upper contour sets stay at a positive distance from their separated feasible counterparts. An application of Lemma 1 to appropriately chosen subsets implies the probability bounds in Theorems 1 and 2.
We begin by laying down some terminology. For a vector $p\in\mathbb{R}^{d}$, $p\neq\bm{0}$, and a constant $b$, we define two half-spaces:
[…]
@@
where the last inequality holds because $p\in\mathbb{R}_{+}^{d}$ and $\omega_{i}\geq\tau\bm{1}$.
Therefore, since $i\in I$ and $g_{i}\in\mathcal{U}_{i}^{(\varepsilon)}(f_{i})$ were arbitrary, by the above inequality $\mathcal{U}_{i}^{(\varepsilon)}(f_{i})-\{\omega_{i}\}\subseteq H^{+}\left(p;\varepsilon\tau\lVert p\rVert_{1}\right)$ for all $i\in I$. Let us define $\mathcal{Q}\coloneq\bigcup_{i\in I}\left(\mathcal{U}_{i}^{(\varepsilon)}(f_{i})-\{\omega_{i}\}\right)$. [^9] The set $\mathcal{Q}$ was originally used by Debreu and Scarf (1963) to prove core convergence, and by Barman and Echenique (2023) to characterize approximate Walrasian equilibria. Then $\mathcal{Q}\subseteq H^{+}\left(p;\varepsilon\tau\lVert p\rVert_{1}\right)$, so for an arbitrary $r>0$, one has
@@
The equality above follows from (5.4), and the last inequality holds because $\min_{p\neq\bm{0}}\lVert p\rVert_{1}/\lVert p\rVert_{2}=1$, namely the minimum of $\lVert p\rVert_{1}/\lVert p\rVert_{2}$ is achieved on the standard unit basis vectors, and is equal to $1$.
Now set $A\coloneq\mathcal{Q}\cap\mathbb{B}(r)$ and $B\coloneq H^{-}(p;0)\cap\mathbb{B}(r)$. By the above inequality $\mathsf{dist}(A,B)\geq\varepsilon\tau$. Since $p$ is a nonzero vector in $\mathbb{R}_{+}^{d}$, the subset $B$ covers at least half of the volume of $\mathbb{B}(r)$. So it must be that $\mathsf{Vol}(A)\leq\mathsf{Vol}(B)$. Therefore, Lemma 1 implies that $\mathsf{Vol}(A)/\mathsf{Vol}(\mathbb{B}(r))\leq\mathrm{e}^{-\varepsilon^{2}\tau^{2}d/8r^{2}}$, namely:
@@
Now suppose $v\in\mathcal{V}^{(\varepsilon)}$. Then, there are $g_{i}\in\mathcal{U}_{i}^{(\varepsilon)}$ for every $i\in I$, such that $v=\sum_{i\in I}g_{i}$ and $(1-\varepsilon)g_{i}\succ_{i}f_{i}$ . For each $i\in I$, it holds that $(1-\varepsilon)g_{i}\in\mathcal{U}_{i}^{(0)}$ and hence $(1-\varepsilon)\sum_{i\in I}g_{i}\in\mathcal{V}^{(0)}$. Consequently, one has $p\cdot(1-\varepsilon)v\geq\lVert p\rVert_{1}$. That in turn implies $v\in H^{+}\left(p;\lVert p\rVert_{1}/(1-\varepsilon)\right)$, and thereby $\mathcal{V}^{(\varepsilon)}\subseteq H^{+}\left(p;\lVert p\rVert_{1}/(1-\varepsilon)\right)$. As a result of this set inclusion, for an arbitrary $r>0$, one obtains that
[…]